OpenStudy (anonymous):

SOLVING SYSTEMS OF EQUATIONS ALGEBRAICALLY y=4x+45 WILL GIVE A MEDAL!!! x=4Y

OpenStudy (anonymous):

y = 4(4y) + 45 y = 16y + 45 -15y = 45 y = -3 x = 4(-3) x = -12 (-12,-3) is your answer :)